OpenStudy (anonymous):

Please help with one Algebra question.? :) 13. If an object is dropped from a height of 144 ft, the function h(t) = –16t² + 144 gives the height of the object after t seconds. When will the object hit the ground? (1 point) 1.5 s 3 s 6 s 9 s

OpenStudy (anonymous):

solve 0= -16*t^2+144. zero here represents the ground.

OpenStudy (anonymous):

use d=1/2at^2+vi

OpenStudy (anonymous):

> -144=-16*t^2 >> 9=t^2 >>>t=3

OpenStudy (anonymous):

Just to make it clearer. What happens when t = 0? We are at our starting point, correct? That means that the exercise picked 0 to be the ground, as our y intercept is 144. That is why we have to solve the quadratic: \( -16t^2 + 144 \). In general, these Physics problems reduce to understanding what is your reference system, i.e., where is the origin or the zero.